OpenStudy (anonymous):

convert these rectangular coordinates to polar coordinates: (-4,3)

OpenStudy (jdoe0001):

\(\large{ \begin{array}{cccllll} (&-4,&\quad 3)\\ &\uparrow &\uparrow \\ &x&y \end{array} \\ \quad \\ r^2=x^2+y^2\implies {\color{brown}{ r}}=\sqrt{x^2+y^2} \\ \quad \\ tan(\theta)=\cfrac{y}{x}\to tan^{-1}[tan(\theta)]=tan^{-1}\left(\cfrac{y}{x}\right)\to {\color{brown}{ \theta}}=tan^{-1}\left(\cfrac{y}{x}\right) \\ \quad \\ (\square ,\square )\implies ({\color{brown}{ r}},{\color{brown}{ \theta}}) }\)

OpenStudy (anonymous):

(5,-36.86) ?